LOW TEMPERATURE FTIR SPECTRA OF METHYL SILANE

\underline{CHANGHONG XIA}, GEOFFREY DUXBURY^{a}, Department of Physics and Applied Physics, Strathclyde University, Glasgow G4 0NG,Scotland,UK; DAVID NEWNHAN, and JOHN BALLARD, Atmospheric Science Division, Space Science Department, Rutherford Appleton Laborator, Chilton, Didcot, OX11 0QX, UK.

The fundamental vibrational bands of methyl silane (CH_{3}SiH_{3}) ν_{1}, ν_{2}, ν_{7}, and ν_{8} have been recorded at 203K and 77K using both conventional absorption cells and also the RAL collisional cooling cell, at resolutions of up to 0.0017cm-1. In the low temperature spectra, the torsional splitting of the rotational transitions, which is obscured in the room temperature spectra, has been resolved and assigned in the majority of cases. Previously missasignment of the ν_{1} Q branch splitting has corrected. Rotation constants of methyl silane and the torsional barrier for the different excited states have been determined

avisiting fellow at JILA